# Integrated population model reveals human and environment driven changes in Baltic ringed seal (Pusa hispida botnica) demography and behavior

### v1

@@ Abstract
Integrated population models (IPMs) are a promising approach to test ecological theories and assess wildlife populations in dynamic and uncertain conditions. By combining multiple data sources into a single, unified model, they enable the parametrization of versatile, mechanistic models that can predict population dynamics in novel circumstances. Here, we present a Bayesian IPM for the ringed seal (Pusa hispida botnica) population inhabiting the Bothnian Bay in the Baltic Sea. Despite the availability of long-term monitoring data, traditional assessment methods have faltered due to dynamic environmental conditions, varying reproductive rates, and the recently re-introduced hunting, thus limiting the quality of information available to managers. We fit our model to census and various demographic, reproductive and harvest data from 1988 to 2023 to provide a comprehensive assessment of past population trends, and predict population response to alternative hunting scenarios. We estimated that 20,000 to 36,000 ringed seals inhabit the Bothnian Bay, and the population is increasing 3% to 6% per year. Reproductive rates have increased since 1988, leading to a substantial increase in the growth rate up until 2015. However, the re-introduction of hunting has since reduced the growth rate, and even minor quota increases are likely to reduce it further. Our results also support the hypothesis that a greater proportion of seals haul-out under lower ice cover circumstances, leading to higher aerial survey counts in such years. In general, our study demonstrates the value of IPMs for monitoring natural populations under changing environments, and supporting science-based management decisions.
@@
integrated population model; state-space model; Bayesian; Baltic; ringed seal; Pusa hispida; demography; wildlife management; environmental change; monitoring;
@@ 1 Introduction
There is, thus, a growing need for versatile, mechanistic population models capable of testing alternative hypotheses on development of natural populations and accommodating uncertainties in population assessments. The integration of such models with long term monitoring data can lead to efficient generation of explanatory and anticipatory predictions (mouquet2015; maris2018; marolla2021). The repeated application of short-term predictions, combined with informed management decisions and continuous monitoring, can facilitate testing and refining model assumptions and accelerate the pace of research - a process that is at the heart of adaptive management (holling1978; lahoz2014; dietze2018).
@@
The demand for mechanistic, data-driven population models was among the key motivations behind the development of integrated population models (IPMs), which combine multiple data sources in a single, unified model to simultaneously infer key demographic parameters and population processes (besbeas2002; Buckland2004; schaub2011; zipkin2018). This is in contrast to traditional approaches where independent empirical estimates for model parameters are typically incorporated into a population projection matrix such as a Leslie matrix (caswell2001). The main advantages of IPMs include their flexibility, their ability to separate process variability from observation error, and their ability to more precisely estimate a larger number of parameters by leveraging synergies between multiple data sources (schaub2011; abadi2010; zipkin2018). Additional information can also be incorporated under the Bayesian framework in the form of prior distributions, which can be determined based on previous empirical work on similar species or expert opinions (Buckland2004). Sufficiently developed IPMs could potentially function as ”digital twins” of their target populations, where the data is continuously updated and the model recalibrated in a way that enables managers to swiftly respond to changing conditions (de2023; trantas2023). Despite the great promise of IPMs as digital twins of wild animal populations, this potential remains unrealized.
@@
To demonstrate the value of IPMs in monitoring and managing animal populations in dynamic and uncertain conditions, we developed a Bayesian IPM for the Baltic ringed seal (Pusa hispida botnica) population inhabiting the Bothnian Bay, the northernmost region of the Baltic Sea. We adopted a state-space formulation for our model (Buckland2004), and parametrized it using count data obtained from aerial transect surveys, demographic data from hunted and by-caught seal samples, hunting records from Finland and Sweden, and assessments of reproductive status in sampled females. The ringed seal population in the Bothnian Bay is a rare example of an ice-dependent pinniped population for which long term monitoring data is available. However, despite the abundance of data, uncertainties stemming from a multitude of changing conditions have precluded assessments of population size and growth for over a decade (HELCOMindicator).
@@
The population size of Baltic ringed seals had plummeted from an estimated 100,000-450,000 in the year 1900 to about 5,000 by the late 1970s, driven largely by unsustainable bounty hunting practices and extremely low reproductive rates caused by organochlorine contamination (helle1976; durant1986; Harkonen1998; Harding1999; Kokko1999; Harkonen2008). Both seal hunting and the use of PCB and DDT were subsequently prohibited, and the Bothnian Bay population has shown promising signs of recovery since the late 1980s. Aerial transect surveys conducted in 1988-2012 suggested an annual growth rate of approximately 5% (Sundqvist2012). However, more recent census estimates have shown unusually high variability and systematic deviations from historical trends, casting uncertainty over the current status of the population (HELCOMindicator).
@@
In the Bothnian Bay, aerial transect surveys are conducted annually around the third week of April, and ringed seals are counted when they are visible while molting on the sea ice after having abandoned their subnivean lairs (Harkonen1992). Results from aerial surveys provide an index of population size and trend under the assumption that approximately the same proportion of the population is hauled-out on ice during the surveys each year (Harkonen1992). It has been speculated, however, that a substantially larger fraction of seals may haul-out when aerial surveys coincide with low ice cover or early ice breakup, resulting in extremely large population counts that are not comparable to typical results (HELCOMindicator). Moreover, previous analyses of survey estimates have often assumed constant population growth, implying constant demographic rates (Sundqvist2012). However, reproductive rates of Baltic ringed seals have likely been improving since the use of PCB and DDT were banned (helle1980; Kauhala2019; HELCOMindicator_reproduction). Unsurprisingly, the population growth rate may have been increasing as well (HELCOMindicator). Thus, we hypothesized that the atypical aerial survey estimates observed over the past decade stem from a combination of improvements in reproductive rates and a higher visibility of seals on ice following mild winters.
@@
The challenges of managing the ringed seal population in the Bothnian Bay are amplified by the recent re-introduction of seal hunting in Finland and Sweden in an attempt to mitigate the rising tension between fisheries and the growing seal population. The effectiveness of hunting as a strategy for alleviating seal-fishery conflicts hinges on achieving a compromise between conservation goals and the interests of coastal fisheries (oksanen2014; cummings2019). This requires reliable assessments of population size, growth rate and demography, an understanding of seal-human interactions, as well as the ability to predict population responses to alternative management decisions. However, due to the ongoing difficulties with monitoring ringed seals in the Bothnian Bay, the impact of recently established hunting on ringed seal demography is unknown.
@@ 2.1 Study population
The Baltic ringed seal is a subspecies of ringed seal endemic to the Baltic Sea (rice1998; but see palo2001). Recognized sub-populations inhabit the Bothnian Bay, the Archipelago Sea, the Gulf of Riga and the Gulf of Finland (Harkonen1998; halkka2017). Our study focused on the Bothnian Bay population, which comprises over 75% of all Baltic ringed seals (Sundqvist2012). The Bothnian Bay is the northernmost region of the Baltic Sea, with a surface area of approximately 37,000 km2 (Figure 1). Since the southern sub-populations in the Archipelago Sea, the Gulf of Finland and the Gulf of Riga are small (Sundqvist2012), and ringed seals show high breeding site fidelity (Kelly2010), we treated the dynamics of the Bothnian Bay population independently of the other sub-populations.
@@
During the foraging period between June and January, Baltic ringed seals spend over 90% of their time in the water, hauling-out on land primarily at night for resting (Oksanen2015). This phase is marked by intense feeding activities, as ringed seals seek to accumulate blubber reserves in preparation for the approaching winter (Oksanen2015; Kauhala2019). Interactions with coastal fisheries are therefore most common between late spring and late fall. During this period, bycatch mortality is most typical for young seals (oksanen2015b; jounela2019).
@@
During the winter, ringed seals may spend up to 20% of their time in subnivean lairs built on sea ice (Kelly2010). Around February-March, a single pup is born inside the snow lair, and nursed for 5-8 weeks (helle1979b; lydersen1993). The snow lair provides shelter to pups against harsh weather conditions and predators. Therefore, pups born without the protection offered by a stable lair are unlikely to survive (smith1975; Ferguson2005; Sundqvist2012). Because whelping and nursing require sufficient ice and snow conditions, adult seals are often associated with stable pack ice during the breeding season, whereas juveniles are more commonly observed in productive foraging areas near the ice edge or in the open water (crawford2012; Oksanen2015). Mating is thought to take place within one month of parturition (mclaren1958; stirling1983), although the implantation of the embryo is delayed until around July (mclaren1958). If seals do not have enough blubber reserves, implantation may not occur (boyd1999).
@@
The molting period begins as the snow cover starts to melt and subnivean lairs collapse in April. Molting ringed seals spend a significant portion (50-80%; Harkonen1992; born2002; Kelly2010) of their time basking on the sea ice in order to increase blood flow to the skin and maintain elevated skin temperatures (feltz1966; thometz2021). Sub-adults and adults often continue molting until the ice melts, typically around late May in the Bothnian Bay (helle1980survey; Harkonen2008), whereas pups typically complete their natal hair molt in lairs (smith1970; lydersen1993).
@@
The molting period also marks the beginning of the ringed seal hunting season in Sweden and Finland, which was re-introduced in 2015 and in 2016, respectively. Some small scale hunting took place earlier, mainly for research purposes. Annual hunting quotas were initially set to about 100 seals in each country, and have gradually been raised to 420 seals in Sweden and 375 seals in Finland as of 2022 (Section 2.2.2). Sweden reduced quotas to 350 seals in 2023, whereas the Finnish quota was kept the same.
@@
Hunting in Sweden is legal between May 1 and January 15, although the actual distribution of hunting activities is bimodal, with distinct spring and autumn seasons (Figure 2). Sweden practices ”protective” hunting, which largely restricts hunting to within 200m of fishing sites where seals have previously caused damage to gear or taken catch. Protective hunting is also allowed within the vicinity of fish farms, fish conservation areas and fish release sites. Hunting in Sweden is therefore likely to be opportunistic.
@@
Hunting in Finland is legal between April 16 and December 31. Finland allows recreational hunting of ringed seals, and hunters are therefore more likely to actively seek out seals. Hunting in Finland almost exclusively takes place on sea ice, and the hunting season effectively ends once sea ice melts (Figure 2). The size of the Finnish hunting bag is influenced by the severity of winter. For example, following mild winters seals tend to congregate in smaller areas of stable ice closer to land, making them more accessible to hunters (Harkonen1992; Harding1999; managementplan2007).
@@ 2.2.1 Aerial surveys
The basking population size of ringed seals during the peak molting season was estimated by the Swedish Museum of Natural History (SMNH) using aerial strip surveys conducted in the Bothnian Bay in late April each year in 1988-2023. The 800 m wide strips were placed evenly over the study area, and the number of observed seals was recorded from an aircraft flying at 90 m altitude (Harkonen1992; Harkonen1998). The proportion of the total ice area covered by the surveys ranged from 13% to 53%, with the exception of 2016, when only 7% of the total ice area was surveyed. The average density of seals over the surveyed transects was extrapolated over the whole ice covered area to estimate the total size of the basking population. The precision of the estimates improves with increasing coverage, although the gain in precision begins to diminish beyond a coverage of about 13% (Harkonen1992).
2.2.2 Hunting bag and quotas
The sizes of the Finnish hunting quotas and bags from 2016-2022 were provided by the Finnish Wildlife Agency. Hunting records, quotas and sampling protocols from Sweden between 2015-2022 were provided by SMNH.
@@ 2.2.3 Demographic and reproductive data
We used only samples of females that were 5+ years old for the post-partum signs, and 4+ years old for a visible fetus (see Section 2.3.1). Samples which were not evaluated for any reproductive sign were excluded, resulting in yearly sample sizes ranging from 0-18 for visible embryos, 0-62 for placental scars and 0-83 for CA (Appendix S1).
From here on, we define the pregnancy rate as the proportion of 4+ year old females carrying a fetus during the fall. In order to account for possible pregnancy losses, postpartum signs of pregnancy were modeled separately. We define the birth rate as the proportion of 4+ year old females that remain pregnant long enough to have a placental scar at the time of parturition, as this is thought to be the closest measurement available to assess the percentage of adult females that produce a live pup. Our definition of birth may include some late-term abortions and stillbirths, which we consider pup mortality.
@@ 2.3.1 Structure of the population dynamics model
We modeled the population dynamics of ringed seals using an age and sex structured model with demographic stochasticity in births and deaths. Because female ringed seals typically begin reproducing at the age of five years (mclaren1958; Lydersen1987; Reimer2019; Kauhala2019), and reproductive senescence is rare (mclaren1958; ellis2018), we included six age classes in our model, grouping 5+ year olds in a single age class. We refer to 0 year old individuals as pups, 1-4 year olds as sub-adults, and 5+ year olds as adults. As hunting quotas do not discriminate between males and females, and the sex ratio of the hunting bag does not necessarily reflect the sex ratio of the population, we included males in our population model but assumed that females were the limiting sex in reproduction.

### v2

@@ Abstract
Integrated population models (IPMs) are a promising approach to test ecological theories and assess wildlife populations in dynamic and uncertain conditions. By combining multiple data sources into a unified model, they enable the parametrization of versatile, mechanistic models that can predict population dynamics in novel circumstances. Here, we present a Bayesian IPM for the ringed seal (Pusa hispida botnica) population inhabiting the Bothnian Bay in the Baltic Sea. Despite the availability of long-term monitoring data, traditional assessment methods have faltered due to dynamic environmental conditions, varying reproductive rates, and the recently re-introduced hunting, thus limiting the quality of information available to managers. We fit our model to census and various demographic, reproductive, and harvest data from 1988 to 2023 to provide a comprehensive assessment of past population trends, and predict population response to alternative hunting scenarios. We estimated that 20,000 to 36,000 ringed seals inhabited the Bothnian Bay in 2024, increasing at a rate of 3% to 6% per year. Reproductive rates have increased since 1988, leading to a substantial increase in the growth rate up until 2015. However, the re-introduction of hunting has since reduced the growth rate, and even minor quota increases are likely to reduce it further. Our results also support the hypothesis that a greater proportion of the population hauls out under lower ice cover circumstances, leading to higher aerial survey results in such years. In general, our study demonstrates the value of IPMs for monitoring wildlife populations under changing environments, and supporting science-based management decisions.
@@
IPM; state-space model; Bayesian; pinniped; environmental change; haulout; long-term monitoring; wildlife management;
@@ 1 Introduction
There is, thus, a growing need for versatile, mechanistic population models capable of testing alternative hypotheses on development of natural populations and accommodating uncertainties in population assessments. The integration of such models with long term monitoring data can lead to efficient generation of explanatory and anticipatory predictions (mouquet2015; maris2018; marolla2021). The repeated application of short-term predictions, combined with informed management decisions and continuous monitoring, can facilitate testing and refining model assumptions and accelerate the pace of research – a process that is at the heart of adaptive management (holling1978; lahoz2014; dietze2018).
@@
The demand for mechanistic, data-driven population models was among the key motivations behind the development of integrated population models (IPMs), which combine multiple data sources in a single, unified model to simultaneously infer key demographic parameters and population processes (besbeas2002; Buckland2004; schaub2011; zipkin2018). This is in contrast to traditional approaches where independent empirical estimates for model parameters are typically incorporated into a population projection matrix such as a Leslie matrix (caswell2001). The main advantages of IPMs include their flexibility, their ability to separate process variability from observation error, and their ability to more precisely estimate a larger number of parameters by leveraging synergies between multiple data sources (abadi2010; schaub2011; zipkin2018). Additional information can also be incorporated under the Bayesian framework in the form of prior distributions, which can be determined based on previous empirical work on similar species or expert opinions (Buckland2004). Sufficiently developed IPMs could potentially function as ”digital twins” of their target populations, where the data is continuously updated and the model recalibrated in a way that enables managers to swiftly respond to changing conditions (de2023; trantas2023; lecarpentier2024). Despite the great promise of IPMs as digital twins of wild animal populations, their potential has yet to be fully realized.
@@
To demonstrate the value of IPMs in monitoring and managing animal populations in dynamic and uncertain conditions, we developed a Bayesian IPM for the Baltic ringed seal (Pusa hispida botnica) population inhabiting the Bothnian Bay, the northernmost region of the Baltic Sea and home to over 75% of all Baltic ringed seals (Harkonen1998; Sundqvist2012; halkka2017). We adopted a state-space formulation for our model (Buckland2004), and parametrized it using count data obtained from aerial transect surveys, demographic data from hunted and by-caught seal samples, hunting records from Finland and Sweden, and assessments of reproductive status in sampled females. The ringed seal population in the Bothnian Bay is a notable example of an ice-dependent pinniped population for which long term monitoring data is available. However, despite the abundance of data, uncertainties stemming from a multitude of changing conditions have precluded assessments of population size and growth for over a decade (HELCOMindicator).
@@
The population size of Baltic ringed seals had plummeted from an estimated 100,000-450,000 in the year 1900 to about 5,000 by the late 1970s, driven largely by unsustainable bounty hunting practices and extremely low reproductive rates caused by organochlorine contamination (helle1976; durant1986; Harkonen1998; Harding1999; Kokko1999; Harkonen2008). Both seal hunting and the use of PCB and DDT were subsequently prohibited, and the Bothnian Bay population has shown signs of recovery since the late 1980s. Aerial transect surveys conducted in 1988-2012 suggested an annual growth rate of approximately 5% (Sundqvist2012). However, more recent census estimates have shown unusually high variability and systematic deviations from historical trends, casting uncertainty over the current status of the population (HELCOMindicator).
@@
In the Bothnian Bay, aerial transect surveys are conducted annually around the third week of April, and ringed seals are counted when they are visible while molting on the sea ice (i.e. ”hauled out”) after having abandoned their subnivean lairs (Harkonen1992). Aerial survey results have traditionally been used as an index of population size and trend, based on the assumption that approximately the same proportion of the population hauls out on ice during the surveys each year (Harkonen1992). It has been speculated, however, that a substantially larger fraction of seals may haul out when aerial surveys coincide with low ice cover or early ice breakup, resulting in extremely large population counts that are not comparable to typical results (HELCOMindicator). Moreover, previous analyses of survey estimates have often assumed constant population growth, implying constant demographic rates (Sundqvist2012). However, reproductive rates of Baltic ringed seals have likely been improving since the use of PCB and DDT were banned (helle1980; Kauhala2019; HELCOMindicator_reproduction). Unsurprisingly, the population growth rate may have been increasing as well (HELCOMindicator). Thus, we hypothesized that the atypical aerial survey estimates observed over the past decade stem from a combination of improvements in reproductive rates and a higher visibility of seals on ice following mild winters.
@@
The challenges of managing the ringed seal population in the Bothnian Bay are amplified by the recent re-introduction of seal hunting in Finland and Sweden in an attempt to mitigate the rising tension between fisheries and the growing seal population. The effectiveness of hunting as a strategy for alleviating seal-fishery conflicts hinges on achieving a compromise between conservation goals and the interests of coastal fisheries (oksanen2014; cummings2019). This requires reliable assessments of population size, growth rate and demography, an understanding of seal-human interactions, as well as the ability to predict population responses to alternative management decisions. However, due to the ongoing difficulties with monitoring ringed seals in the Bothnian Bay, the impact of hunting on ringed seal demography is unknown.
@@ 2.1 Study population
The Baltic ringed seal is a subspecies of ringed seal endemic to the Baltic Sea (rice1998; but see palo2001). Recognized sub-populations inhabit the Bothnian Bay, the Archipelago Sea, the Gulf of Riga and the Gulf of Finland (Harkonen1998; halkka2017). Our study focused on the Bothnian Bay population, which comprises over 75% of all Baltic ringed seals (Sundqvist2012). The Bothnian Bay is the northernmost region of the Baltic Sea, with a surface area of approximately 37,000 km2 (Figure 1). Since the southern sub-populations in the Archipelago Sea, the Gulf of Finland and the Gulf of Riga are small (Sundqvist2012), and ringed seals show high breeding site fidelity (Harkonen2008; Kelly2010), we treated the dynamics of the Bothnian Bay population independently of the other sub-populations.
@@
During the foraging period between June and January, Baltic ringed seals spend over 90% of their time in the water, hauling out on land primarily at night for resting (Oksanen2015). This phase is marked by intense feeding activities, as ringed seals seek to accumulate blubber reserves in preparation for the approaching winter (Oksanen2015; Kauhala2019). Interactions with coastal fisheries are therefore most common between late spring and late fall. During this period, bycatch mortality is most typical for young ringed seals (oksanen2015b; jounela2019).
@@
During the winter, ringed seals spend most of their time out of the water in subnivean lairs built on sea ice (Kelly2010). Around February-March, a single pup is born inside the snow lair, and nursed for 5-8 weeks (helle1979b; lydersen1993). The snow lair provides shelter to pups against harsh weather conditions and predators. Therefore, pups born without the protection offered by a stable lair are unlikely to survive (smith1975; Ferguson2005; Sundqvist2012). Because whelping and nursing require sufficient ice and snow conditions, adult seals are often associated with stable pack ice during the breeding season, whereas juveniles are more commonly observed in productive foraging areas near the ice edge or in the open water (crawford2012; Oksanen2015). Mating is thought to take place within one month of parturition (mclaren1958; stirling1983), although the implantation of the embryo is delayed until around July (mclaren1958). If females do not have enough blubber reserves, implantation may not occur (boyd1999).
@@
The molting period begins as the snow cover starts to melt and subnivean lairs collapse, typically in April. Molting ringed seals spend a significant portion (50-80%; Harkonen1992; born2002; Kelly2010) of their time basking on the sea ice in order to increase blood flow to the skin and maintain elevated skin temperatures (feltz1966; thometz2021). Sub-adults and adults often continue molting until the ice melts, typically around late May in the Bothnian Bay (helle1980survey; Harkonen2008), whereas pups typically complete their natal hair molt in lairs (smith1970; lydersen1993).
@@
The molting period also marks the beginning of the ringed seal hunting season in Sweden and Finland, which was re-introduced in 2015 and in 2016, respectively. Some small scale hunting took place earlier, mainly for research purposes (nyman2003; routti2009). Annual hunting quotas were initially set to about 100 seals in each country, and have gradually been raised to 420 seals in Sweden and 375 seals in Finland as of 2022 (Section 2.2.2). Sweden reduced quotas to 350 seals in 2023, whereas the Finnish quota was kept the same.
@@
Hunting of ringed seals in Sweden is legal between May 1 and January 15, although the actual distribution of hunting activities is bimodal, with distinct spring and autumn seasons (Figure 2). Sweden practices ”protective” hunting, which largely restricts hunting to within 200m of fishing sites where seals have previously caused damage to gear or taken catch. Protective hunting is also allowed within the vicinity of fish farms, fish conservation areas and fish release sites. Hunting in Sweden is therefore likely to be opportunistic.
@@
Hunting of ringed seals in Finland is legal between April 16 and December 31. Finland allows recreational hunting of ringed seals, and hunters are therefore more likely to actively seek out seals. Hunting in Finland almost exclusively takes place on sea ice, and the hunting season effectively ends once sea ice melts (Figure 2). The total harvest in Finland is influenced by the severity of winter. For example, following mild winters seals tend to congregate in smaller areas of stable ice closer to land, making them more accessible to hunters (Harkonen1992; Harding1999; managementplan2007).
@@ 2.2.1 Aerial surveys
The basking population size of ringed seals during the peak molting season was estimated by the Swedish Museum of Natural History (SMNH) using aerial strip surveys conducted in the Bothnian Bay in mid to late April each year between 1988 and 2023. The 800 m wide strips were placed evenly over the study area, and the number of observed seals was recorded from an aircraft flying at 90 m altitude (Harkonen1992; Harkonen1998). The proportion of the total ice area covered by the surveys ranged from 13% to 53%, with the exception of 2016, when only 7% of the total ice area was surveyed. The average density of seals over the surveyed transects was extrapolated over the whole ice covered area to estimate the total size of the basking population, which is thought to account for 50-80% of the overall population (Harkonen1992; born2002; Kelly2010). The precision of the estimates improves with increasing coverage, although the gain in precision begins to diminish beyond a coverage of about 13% (Harkonen1992).
2.2.2 Harvest totals and quotas
Finnish hunting quotas and harvest totals from 2016-2022 were provided by the Finnish Wildlife Agency. Hunting records, quotas and sampling protocols from Sweden between 2015-2022 were provided by SMNH.
@@ 2.2.3 Demographic and reproductive data
We used only samples of females that were 5+ years old for the post-partum signs, and 4+ years old for a visible fetus (see Section 2.3.1). Samples which were not evaluated for any reproductive sign were excluded, resulting in yearly sample sizes ranging from 0-18 for visible embryos, 0-62 for placental scars and 0-83 for CA (Supplement 1).
From here on, we define the pregnancy rate as the proportion of 4+ year old females carrying a fetus during the fall. In order to account for possible pregnancy losses, postpartum signs of pregnancy were modeled separately. We define the birth rate as the proportion of 4+ year old females that remain pregnant long enough to have a placental scar at the time of parturition, as this is thought to be the closest measurement available to assess the percentage of adult females that produce a live pup. Our definition of birth may include some late-term abortions and stillbirths, which are potentially common in pinnipeds (mckenzie2005; stenson2016) and treated as pup mortality in our analysis.
@@ 2.3.1 Structure of the population dynamics model
We modeled the population dynamics of ringed seals using an age and sex structured model with demographic stochasticity in births and deaths. Because female ringed seals typically begin reproducing at the age of five years (mclaren1958; Lydersen1987; Kauhala2019; Reimer2019), and reproductive senescence is rare (mclaren1958; ellis2018), we included six age classes in our model, grouping 5+ year olds in a single age class. We refer to 0 year old individuals as pups, 1-4 year olds as sub-adults, and 5+ year olds as adults. Note that while 4 year old sub-adults can become pregnant in our model, they do not give birth until reaching adulthood at age 5. As hunting quotas do not discriminate between males and females, and the sex ratio of the harvests do not necessarily reflect the sex ratio of the population, we included males in our population model but assumed that females were the limiting sex in reproduction.